Grade 11Mechanics

A force of 9N is applied to pull the block of mass 4 kg with a string of mass 0.5 kg.Find the force of reaction exerted by block on the rope.

ranu
9 Years agoGrade 11

1 Answer

Shaswata Biswas
ApprovedApproved Tutor Answer9 Years ago
Let mass of rope = mr = 0.5 kg, mass of block = mb = 4 kg, force applied = F = 9 N
Let the tension in the string be T
Common acceleration of the system, ac = F÷(mr+mb) = 9/4.5 = 2 m/s2
For the block to be moving with this acceleration, tension, T = mb.ac = 4*2 = 8 N.
THANKS